# On the accuracy of posterior recovery with neural network emulators

### v1

@@ Abstract
pubyear: 2025pagerange: On the accuracy of posterior recovery with neural network emulatorsC
@@ 1 Introduction
Emulators have been developed in a variety of fields including SED fitting (Alsing et al., 2020), CMB primary and secondary studies (Spurio Mancini et al., 2022) and 21-cm cosmology (Cohen et al., 2020; Bevins et al., 2021b; Bye et al., 2022; Breitman et al., 2024; Dorigo Jones et al., 2024). They are trained on sets of example simulations from numerical or semi-numerical codes to translate between physical parameters such as the Hubble constant or star formation efficiency and return various summary statistics such as a power spectrum. Often when designing emulators, researchers either try to make the emulators as accurate as possible or alternatively use ad hoc rules of thumb to define minimum requirements for the accuracy. For example, when designing the 21-cm emulator globalemu the authors suggested that the emulator should have an average accuracy $\leq 10\%$ of the expected noise in a 21-cm experiment (Bevins et al., 2021b) so that the contribution to the uncertainty from the emulator in any analysis is subdominant to the uncertainty introduced by the instrument.
@@
Dorigo Jones et al. (2023) [hereafter DJ23] began to discuss this topic within the field of 21-cm cosmology. Researchers in this field aim to observe the evolution of the 21-cm signal from the spin-flip transition of neutral hydrogen. The signal traces the formation of the first stars and galaxies through to the moment at which the Universe transitioned from being predominantly neutral to ionized. The signal can be observed in the radio band as a sky-averaged signature with a single radiometer, or alternatively as a power spectrum with an interferometer. Although no current confirmed detections of either the power spectrum or the sky-averaged signal exist, there are a number of upper limits on both and emulators have been used extensively to analyse these data sets (e.g. Bevins et al., 2022c, b; Abdurashidova et al., 2022; HERA Collaboration et al., 2023; Monsalve et al., 2019).
@@
In section 2 we define the upper limit on the KL divergence between the true and emulated posterior as a function of the noise in the data and emulator accuracy. We then demonstrate this with ARES and globalemu in section 3. We conclude in section 4. Our code is publicly available at https://github.com/htjb/validating_posteriors.
@@ 2 Measuring information loss
If an emulator is used in a Bayesian inference pipeline, then the recovered posterior $P_{\epsilon}(\theta|D,M_{\epsilon})$ will differ from the true posterior $P(\theta|D,M)$. For the emulator to be useful then the difference between these two posteriors should be negligible. A natural way to quantify the difference between these two probability distributions is the Kullback-Leibler divergence $\mathcal{D}_{\mathrm{KL}}$. Generally the $\mathcal{D}_{\mathrm{KL}}$ between two distributions quantifies the information loss or gain when moving from one to the other. Here the $\mathcal{D}_{\mathrm{KL}}$ quantifies the incorrect information inferred about the parameter space, in natural bits, when using the emulator to recover the posterior. Ideally, we would like the $\mathcal{D}_{\mathrm{KL}}$ between these two posterior distributions to be as close to zero as possible and $<1$ so that no significant false information is inferred.
@@ 2.1 The general case
As above, we will refer to these posteriors as $P$ and $P_{\epsilon}$ for brevity.
We can calculate the KL-divergence of this change in posterior as
@@
Substituting in eq. 3 into the $\log$,
[…]
and using a Taylor expansion of the exponential in eq. 6 we can write the KL divergence as
In the limit that $\delta\log L$ is small, we can consider the low order expansion of $\mathcal{D}_{\mathrm{KL}}$. Noting that $\lambda_{n}=O(\delta\log L{}^{n})$ and assuming $\left|\sum_{i=1}^{\infty}\lambda_{i}\right|<1$ then
$\begin{split}\mathcal{D}_{\mathrm{KL}}&\approx-\lambda_{1}+\left(\sum_{i=1}^{\infty}\lambda_{i}\right)-\frac{1}{2}\left(\sum_{i=1}^{\infty}\lambda_{i}\right)^{2}+\frac{1}{3}\left(\sum_{i=1}^{\infty}\lambda_{i}\right)^{3}-\ldots\\ &=\lambda_{2}-\frac{1}{2}\lambda_{1}^{2}+O\left(\delta\log L{}^{3}\right),\end{split}$ (9)
@@ 2.2 The linear model case
where $\Sigma$ is the data covariance matrix. Substituting in the linear model in eq. 11 and accounting for our assumed prior we get the following Gaussian posterior distribution
@@
with the equivalents for the emulator analogously found using eq. 13 as
@@
These are all standard results (Petersen & Pedersen, 2008).
@@ 2.3 White noise and E=0
$\mathcal{D}_{\mathrm{KL}}\leq\frac{1}{2}\frac{1}{\sigma^{2}}\max(\lambda_{\rm n})\sum_{j}f_{j}^{2}=\frac{1}{2}\frac{1}{\sigma^{2}}\max(\lambda_{\rm n})||f||^{2}.$ (30)
But, since $U^{-1}$ is a rotation matrix it leaves the 2-norm of the vector it acts on unchanged, hence $||f||=||\epsilon||$ and so
$\mathcal{D}_{\mathrm{KL}}\leq\frac{1}{2}\frac{1}{\sigma^{2}}\max(\lambda_{\rm n})||\epsilon||^{2}.$ (31)
@@
where RMSE is the root mean squared error across a test data set for the emulator. Therefore, under the assumptions of Gaussian noise and a linear model for the data, we can say that for less than 1 bit of difference between $P$ and $P_{\epsilon}$ then
@@ 2.4 Validity of the assumptions made
where $\nu$ is frequency of the observations, $T_{A}$ is the antenna temperature, $\Delta\nu$ is the channel width of the data and $\tau$ is the integration time for the observations. In a narrow bandwidth $\sigma(\nu)$ is approximately constant and most studies currently assume a constant level of noise across the observed frequency range. However, $\sigma$ can be replaced in eq. 34 with $\mathrm{min}(\sigma(\nu))$.
Scheutwinkel et al. (2023) explored the use of other likelihood functions in the field of 21-cm cosmology. Since we are assuming the prior is uniform, then the posterior in eq. 15 is approximately equivalent to the likelihood function. If an analytic expression exists for the $\mathcal{D}_{\mathrm{KL}}$ between the two approximate posterior distributions, equivalent to eq. 6, then similar arguments to those proposed here can be followed to arrive at an upper bound on the KL divergence between $P$ and $P_{\epsilon}$.
@@ 3 An example in 21-cm Cosmology
Numerical, Semi-numerical, 1D radiative transfer codes and analytic models are all used to model the evolution of the 21-cm signal and parameterize its dependence on the properties of the first stars and galaxies. Numerical hydro simulations like C2-Ray while detailed are extremely computationally expensive (Mellema et al., 2006). 1D radiative transfer codes like ARES Mirocha et al. (2012); Mirocha (2014) and Zeus21 (Muñoz, 2023) offer a computationally cheaper alternative, but are not as detailed as numerical models. Semi-numerical codes like 21cmSPACE and 21cmFAST can generally be evaluated in a few hours and offer a relatively cheap compromise somewhere between full hydro simulations and 1D codes, allowing for a wider range of physical processes to be modelled.
@@
Several interferometers have placed upper limits on the magnitude of the power spectrum and in 2018 a tentative detection of the sky-averaged signal was made by the EDGES collaboration (Bowman et al., 2018) although this was recently disputed by observations from the SARAS3 instrument (Singh et al., 2022). Previous concerns have also been raised about the cosmological origins of the EDGES signal (Hills et al., 2018; Sims & Pober, 2020; Singh & Subrahmanyan, 2019; Bradley et al., 2019; Bevins et al., 2021a).
A number of works (e.g. Monsalve et al., 2019; Abdurashidova et al., 2022; Singh et al., 2018; HERA Collaboration et al., 2023; Bevins et al., 2022b, c; Pochinda et al., 2024; Gessey-Jones et al., 2024; Bevins et al., 2024) have used upper limits on the magnitude of the sky-averaged or global 21-cm signal and the power spectrum to constrain the properties of the first stars and galaxies. Since 1D radiative transfer codes can be evaluated on the order of seconds, they can be used in Bayesian inference pipelines. However, to make use of semi-numerical and numerical codes, researchers rely on neural network emulators because it is computationally infeasible to run inference directly on the semi-numerical simulations which take of order hours to evaluate per parameter set. Several emulator frameworks of the 21-cm signal exist, such as 21cmVAE (Bye et al., 2022), 21cmGEM (Cohen et al., 2020) and 21cmEMU (Breitman et al., 2024). In this work, we compare the posteriors recovered when using the 1D radiative transfer code ARES and an emulator of ARES built with globalemu to fit mock data of varying noise levels.
@@ 3.2 Previous work
We now summarise the analysis presented in. First, a data set of 21-cm global signal simulations were generated using the ARES code. globalemu emulators were then trained on these simulations. Using these emulators as well as the ARES code directly, inference was performed on synthetic global 21-cm signal data with various noise levels, both in isolation and jointly with UV luminosity function (UVLF). The resulting 1D posteriors were then compared using two different metrics
@@
globalemu includes several physically motivated preprocessing steps to improve the accuracy of trained emulators. These were switched off in DJ23. The paper claims that this makes the emulators more accurate, in contrast to the conclusions in the original globalemu paper. In section 3.4, we show that more accurate emulators can be trained with the preprocessing steps on using the training data from DJ23.
@@ 3.3 ARES setup
We use the same prior ranges as in DJ23 on our parameters and the same fiducial values for our ground truth model (see table 1) to which we add Gaussian distributed noise with a standard deviation of 5, 25, 50 and 250 mK in our experiments.
@@
In table 2 we report the accuracy of the emulation when switching on and off the preprocessing steps discussed above. In the original globalemu paper, the accuracy of the emulator was assessed using an unseen test data set and over fitting was checked for after training by comparing the distribution of the error in predicted signals for the test and training data. More recent iterations of the globalemu code implemented a form of early stopping algorithm using the test data set provided to the emulator. In several works, (e.g. Bevins et al., 2022c, b; Pochinda et al., 2024; Gessey-Jones et al., 2024) including DJ23 the same test data set used for early stopping was used to assess the accuracy of the emulator. In practice, this is not a fair representation of the performance of the emulator because the emulator has been fine-tuned to perform well on that specific data set. In this work, we generate a new set of simulations with ARES to measure the accuracy of the network.
We train six different versions of the emulator for each combination of preprocessing steps. The accuracy for each was then evaluated over a test data set comprising 2000 models over the band $z=6-55$ and the mean Root Mean Squared Error (RMSE), the 95th percentile error and the worst emulation error were calculated. For each combination of preprocessing steps, we then took an average of the recorded mean RMSE values along with a corresponding standard deviation to assess the relative performance of the different emulator set-ups across the stochastic initialisation of the network weights. We did the same for the 95th percentile and worst RMSE values and these are reported in table 2. The emulators are trained on 24,000 simulations covering the prior range outlined in DJ23 and early stopping is performed on a different validation set of 2000 signals. We find that the best mean and 95th percentile values are recovered when both the AFB subtraction and the resampling are switched on.
@@
95th Percentile $3.96\pm 0.72$ $3.98\pm 0.57$ $4.59\pm 0.53$ $4.48\pm 0.33$ 3.14
Worst $31.04\pm 7.09$ $31.58\pm 9.48$ $33.91\pm 6.13$ $33.89\pm 4.70$ 25.97 18.5
Table 2: A comparison of emulator accuracy over the band $z=6-55$ as a function of preprocessing steps. We train six different emulators and assess the accuracy of each on 2000 previously unseen signal models for each combination of preprocessing steps. We report the average values over the six training runs and use the standard deviation to show the stochastic variation in the accuracy. While there is some overlap in the performance for the different set-ups, we find that it is possible to train more accurate emulators with the two preprocessing steps switched on, in agreement with the findings in Bevins et al. (2021b). The most accurate of the emulators that we trained is the one that we use throughout the rest of the paper, and has a mean error of 0.99 mK over the test data set. The values for the DJ23 emulator were taken from their paper. 95 % of the training data have an RMSE lower or equal to the $95^{\mathrm{th}}$ percentile.
We use the same network architecture as DJ23, namely three hidden layers of 32 nodes each, however as previously noted the paper finds a better performance with the AFB subtraction and resampling switched off. We report in table 2 the accuracy of DJ23 emulator for ease of comparison.
@@
In the analysis that follows, we use the best performing emulator trained with AFB and resampling switched on. The accuracy of the emulator as a function of redshift is shown in fig. 1 evaluated on the unseen test data.

### v2

@@ Abstract
pubyear: 2026pagerange: On the accuracy of posterior recovery with neural network emulatorsC
@@ 1 Introduction
Neural Network emulators have been developed in a variety of fields including SED fitting (e.g. Alsing et al., 2020; Mathews et al., 2023, 2025), CMB primary and secondary studies (e.g. Auld et al., 2007; Agarwal et al., 2014; Manrique-Yus and Sellentin, 2020; Albers et al., 2019; Aricò et al., 2021; Mootoovaloo et al., 2022; Spurio Mancini et al., 2022; Bonici et al., 2024; Bolliet et al., 2024; Günther, 2023; Günther et al., 2025), studies of effective field theory of large scale structure (e.g. Bonici et al., 2025), 21-cm cosmology (e.g. Cohen et al., 2020; Bevins et al., 2021b; Bye et al., 2022; Breitman et al., 2024; Dorigo Jones et al., 2024) and others (e.g. El Gammal et al., 2023). They are trained on sets of example simulations from numerical or semi-numerical codes to translate between physical parameters such as the Hubble constant or star formation efficiency and return various summary statistics such as a power spectrum. Often when designing emulators, researchers either try to make the emulators as accurate as possible or alternatively use ad hoc rules of thumb to define minimum requirements for the accuracy. For example, when designing the 21-cm emulator globalemu the authors suggested that the emulator should have an average accuracy $\leq 10\%$ of the expected noise in a 21-cm experiment (Bevins et al., 2021b) so that the contribution to the uncertainty from the emulator in any analysis is subdominant to the uncertainty introduced by the instrument.
@@
Dorigo Jones et al. (2023) [hereafter DJ23] began to discuss this topic within the field of 21-cm cosmology. Researchers in this field aim to observe the evolution of the 21-cm signal from the spin-flip transition of neutral hydrogen. The signal traces the formation of the first stars and galaxies through to the moment at which the Universe transitioned from being predominantly neutral to ionized. The signal can be observed in the radio band as a sky-averaged signature with a single radiometer, or alternatively as a power spectrum with an interferometer. Although no current confirmed detections of either the power spectrum or the sky-averaged signal exist, there are a number of upper limits on both and emulators have been used extensively to analyse these data sets (e.g. Bevins et al., 2022a, b; Abdurashidova et al., 2022; HERA Collaboration et al., 2023; Monsalve et al., 2019).
@@
In section˜2 we define the upper limit on the KL divergence between the true and emulated posterior as a function of the noise in the data and emulator accuracy. We then demonstrate this with ARES and globalemu in section˜3. We conclude in section˜4. Our code is publicly available at https://github.com/htjb/validating_posteriors.
@@ 2 Measuring information loss
If an emulator is used in a Bayesian inference pipeline, then the recovered posterior $P_{\epsilon}(\theta|D,M_{\epsilon})$ will differ from the true posterior $P(\theta|D,M)$. For the emulator to be useful then the difference between these two posteriors should be negligible. A natural way to quantify the difference between these two probability distributions is the Kullback-Leibler divergence $\mathcal{D}_{\mathrm{KL}}$. Generally the $\mathcal{D}_{\mathrm{KL}}$ between two distributions quantifies the information loss or gain when moving from one to the other. Here the $\mathcal{D}_{\mathrm{KL}}$ quantifies the incorrect information inferred about the parameter space, in natural bits, when using the emulator to recover the posterior. Ideally, we would like the $\mathcal{D}_{\mathrm{KL}}$ between these two posterior distributions to be as close to zero as possible and <1<1 so that no significant false information is inferred.
@@ 2.1 The general case
$P(\theta|D,\mathcal{M})=\frac{L\pi}{\int L\pi d\theta}\rightarrow P_{\epsilon}(\theta|D,\mathcal{M}_{\epsilon})=\frac{L\pi\exp(\delta\log L)}{\int L\pi\exp(\delta\log L)d\theta},$ (3)
where $\pi=P(\theta|M)$ is the prior distribution on the free parameters in the model. We will refer to these posteriors as $P$ and $P_{\epsilon}$ for brevity.
We can calculate the KL-divergence of this change in posterior as
$\mathcal{D}_{\mathrm{KL}}=\int P\log\left(\frac{P}{P_{\epsilon}}\right)d\theta.$ (4)
Substituting in eq.˜3 into the $\log$,
$\begin{split}\mathcal{D}_{\mathrm{KL}}&=\int P\log\left(\exp(-\delta\log L)\frac{\int L\pi\exp(\delta\log L)d\theta}{\int L\pi d\theta}\right)d\theta\\ &=-\int P\delta\log Ld\theta+\int P\log\left(\frac{\int L\pi\exp(\delta\log L)d\theta}{\int L\pi d\theta}\right)d\theta\\ &=-\langle\delta\log L\rangle_{P}+\int P\log\left(\int P\exp(\delta\log L)d\theta\right)d\theta.\end{split}$ (5)
$\begin{split}\mathcal{D}_{\mathrm{KL}}&=-\langle\delta\log L{}\rangle_{P}+\log\left(\int P\exp(\delta\log L)d\theta\right)\int Pd\theta\\ &=-\langle\delta\log L{}\rangle_{P}+\log\left(\int P\exp(\delta\log L)d\theta\right)\\ &=\log\left(\frac{\int P\exp(\delta\log L)d\theta}{\exp\left(\int P\delta\log L{}d\theta\right)}\right).\end{split}$ (6)
$\lambda_{n}\equiv\frac{1}{n!}\int P\left(\delta\log L\right)^{n}d\theta$ (7)
and using a Taylor expansion of the exponential in eq.˜6 we can write the KL divergence as
@@ 2.2 The linear model case
where $\Sigma$ is the data covariance matrix. Substituting in the linear model in eq.˜11 and accounting for our assumed prior we get the following Gaussian posterior distribution
@@
with the equivalents for the emulator analogously found using eq.˜13 as
@@
These are all standard results (Petersen and Pedersen, 2008).
@@ 2.3 White noise and E=0
[…]
@@
where RMSE is the root mean squared error across a test data set for the emulator. Therefore, under the assumptions of Gaussian noise and a linear model for the data, we can say that for less than 1 bit of difference [^1] Intuitively, 1 bit of information is the amount gained when flipping a fair coin, since there are two equally likely outcomes of this process. Bits can be thought of as the number of yes/no questions that need to be asked to learn the true answer. Note KL divergence is often calculated in nat bits (base $e$) however this is trivially related to the base 2 KL divergence by $\mathcal{D}_{\mathrm{KL}}^{\mathrm{nat}}=\mathcal{D}_{\mathrm{KL}}^{\mathrm{bits}}\log(2)$. between $P$ and $P_{\epsilon}$ then
@@ 2.4 Validity of the assumptions made
where $\nu$ is frequency of the observations, $T_{A}$ is the antenna temperature, $\Delta\nu$ is the channel width of the data and $\tau$ is the integration time for the observations. In a narrow bandwidth $\sigma(\nu)$ is approximately constant and most studies currently assume a constant level of noise across the observed frequency range. However, $\sigma$ can be replaced in eq.˜34 with $\mathrm{min}(\sigma(\nu))$.
Scheutwinkel et al. (2023) explored the use of other likelihood functions in the field of 21-cm cosmology. Since we are assuming the prior is uniform, then the posterior in eq.˜15 is approximately equivalent to the likelihood function. If an analytic expression exists for the $\mathcal{D}_{\mathrm{KL}}$ between the two approximate posterior distributions, equivalent to eq.˜6, then similar arguments to those proposed here can be followed to arrive at an upper bound on the KL divergence between $P$ and $P_{\epsilon}$.
We note that in higher dimensional spaces, the assumption that the model is linear around the peak of the posterior can become less valid because the posteriors are more complex (increased curvature and/or multimodal structure). Assuming linearity does hold around the peak of the posterior, most of the posterior mass may lie far from the peak, meaning that the KL divergence is dominated by regions in which linearity fails. Although, the limit is not directly dependent on the number of parameters it is dependent on the complexity of the posterior, which is a function of the dimensionality of the model. A full analysis of this behaviour is beyond the scope of this paper and left for future work.
@@ 3 An example in 21-cm Cosmology
Numerical, Semi-numerical, 1D radiative transfer codes and analytic models are all used to model the evolution of the 21-cm signal and parameterize its dependence on the properties of the first stars and galaxies. Numerical hydro simulations like C2-Ray while detailed are extremely computationally expensive (Mellema et al., 2006). 1D radiative transfer codes like ARES Mirocha et al. (2012); Mirocha (2014), Zeus21 (Muñoz, 2023) and ECHO21 (Mittal, 2025) offer a computationally cheaper alternative, but are not as detailed as numerical models. Semi-numerical codes like 21cmSPACE and 21cmFAST can generally be evaluated in a few hours and offer a relatively cheap compromise somewhere between full hydro simulations and 1D codes, allowing for a wider range of physical processes to be modelled.
@@
Several interferometers have placed upper limits on the magnitude of the power spectrum and in 2018 a tentative detection of the sky-averaged signal was made by the EDGES collaboration (Bowman et al., 2018) although this was recently disputed by observations from the SARAS3 instrument (Singh et al., 2022). Previous concerns have also been raised about the cosmological origins of the EDGES signal (Hills et al., 2018; Sims and Pober, 2020; Singh and Subrahmanyan, 2019; Bradley et al., 2019; Bevins et al., 2021a).
A number of works (e.g. Monsalve et al., 2019; Abdurashidova et al., 2022; Singh et al., 2018; HERA Collaboration et al., 2023; Bevins et al., 2022b, a; Pochinda et al., 2024; Gessey-Jones et al., 2024; Bevins et al., 2024) have used upper limits on the magnitude of the sky-averaged or global 21-cm signal and the power spectrum to constrain the properties of the first stars and galaxies. Since 1D radiative transfer codes can be evaluated on the order of seconds, they can be used in Bayesian inference pipelines. However, to make use of semi-numerical and numerical codes, researchers rely on neural network emulators because it is computationally infeasible to run inference directly on the semi-numerical simulations which take of order hours to evaluate per parameter set. Several emulator frameworks of the 21-cm signal exist, such as 21cmVAE (Bye et al., 2022), 21cmGEM (Cohen et al., 2020) and 21cmEMU (Breitman et al., 2024). In this work, we compare the posteriors recovered when using the 1D radiative transfer code ARES and an emulator of ARES built with globalemu to fit mock data of varying noise levels.
@@ 3.2 Previous work
We now summarise the analysis presented in DJ23. First, a data set of 21-cm global signal simulations were generated using the ARES code. globalemu emulators were then trained on these simulations. Using these emulators as well as the ARES code directly, inference was performed on synthetic global 21-cm signal data with various noise levels, both in isolation and jointly with UV luminosity function (UVLF). The resulting 1D posteriors were then compared using two different metrics
@@
globalemu includes several physically motivated preprocessing steps to improve the accuracy of trained emulators. These were switched off in DJ23. The paper claims that this makes the emulators more accurate, in contrast to the conclusions in the original globalemu paper. In section˜3.4, we show that more accurate emulators can be trained with the preprocessing steps on using the training data from DJ23.
@@ 3.3 ARES setup
We use the same prior ranges as in DJ23 on our parameters and the same fiducial values for our ground truth model (see table˜1) to which we add Gaussian distributed noise with a standard deviation of 5, 25, 50 and 250 mK in our experiments.
@@
In table˜2 we report the accuracy of the emulation when switching on and off the preprocessing steps discussed above. In the original globalemu paper, the accuracy of the emulator was assessed using an unseen test data set and over fitting was checked for after training by comparing the distribution of the error in predicted signals for the test and training data. More recent iterations of the globalemu code implemented a form of early stopping algorithm using the test data set provided to the emulator. In several works, (e.g. Bevins et al., 2022a, b; Pochinda et al., 2024; Gessey-Jones et al., 2024) including DJ23 the same test data set used for early stopping was used to assess the accuracy of the emulator. In practice, this is not a fair representation of the performance of the emulator because the emulator has been fine-tuned to perform well on that specific data set. In this work, we generate a new set of simulations with ARES to measure the accuracy of the network.
We train six different versions of the emulator for each combination of preprocessing steps. The accuracy for each was then evaluated over a test data set comprising 2000 models over the band z=655z=6-55 and the mean Root Mean Squared Error (RMSE), the 95th percentile error and the worst emulation error were calculated. For each combination of preprocessing steps, we then took an average of the recorded mean RMSE values along with a corresponding standard deviation to assess the relative performance of the different emulator set-ups across the stochastic initialisation of the network weights. We did the same for the 95th percentile and worst RMSE values and these are reported in table˜2. The emulators are trained on 24,000 simulations covering the prior range outlined in DJ23 and early stopping is performed on a different validation set of 2000 signals. We find that the best mean and 95th percentile values are recovered when both the AFB subtraction and the resampling are switched on.
@@
95th Percentile $3.96\pm 0.72$ $3.98\pm 0.57$ $59\pm 0.53 4.4$$48\pm 0.33$ 3.14
Worst $31.04\pm 7.09$ $1.58\pm 9.48 33$$91\pm 6.13 33.$$3.89\pm 4.70$ 25.97 18.5
Table 2: A comparison of emulator accuracy over the band z=655z=6-55 as a function of preprocessing steps. We train six different emulators and assess the accuracy of each on 2000 previously unseen signal models for each combination of preprocessing steps. We report the average values over the six training runs and use the standard deviation to show the stochastic variation in the accuracy. While there is some overlap in the performance for the different set-ups, we find that it is possible to train more accurate emulators with the two preprocessing steps switched on, in agreement with the findings in Bevins et al. (2021b). The most accurate of the emulators that we trained is the one that we use throughout the rest of the paper, and has a mean error of 0.99 mK over the test data set. The values for the DJ23 emulator were taken from their paper. 95 % of the training data have an RMSE lower or equal to the $95^{\mathrm{th}}$ percentile.
We use the same network architecture as DJ23, namely three hidden layers of 32 nodes each, however as previously noted the paper finds a better performance with the AFB subtraction and resampling switched off. We report in table˜2 the accuracy of DJ23 emulator for ease of comparison.
@@
In the analysis that follows, we use the best performing emulator trained with AFB and resampling switched on. The accuracy of the emulator as a function of redshift is shown in fig.˜1 evaluated on the unseen test data.